# Quantum Fisher information from tensor network integration of Lyapunov equation

### v1

@@ Abstract
Although the role of QFI and SLD is crucial, their evaluation is notoriously difficult because of their nonlinear mathematical form. In general, calculating the QFI requires matrix inversions, which become significant numerical bottlenecks. As a result, considerable effort has been devoted to developing methods for evaluating or bounding the QFI, paralleling advances in quantum metrology. The conventional approach, which relies heavily on eigendecomposition, faces a critical limitation: the Hilbert space dimension grows exponentially with the number of subsystems, making computations intractable for large, composite systems. To address these challenges, alternative formulations that avoid explicit diagonalization have been proposed [13, 14], although they still require certain matrix inversions. Other practical approaches provide QFI bounds using auxiliary measures such as the truncated QFI [15, 16], purity loss [17, 18], and the Krylov shadow method [19]. Recent experimental advances have used noisy intermediate-scale quantum (NISQ) devices to estimate these bounds [15, 16], enabling scalable implementation on physical quantum platforms. In recent decades, tensor network (TN) [20, 21] methods have been shown to provide efficient approximations for relevant classes of states [22, 23, 21], e.g., ground states of gapped Hamiltonians and thermal states of a local Hamiltonian [24, 25]. Among them, the one-dimensional ansatz of the matrix product state (MPS) (also known as the matrix product operator (MPO) for quantum operators) provides a compact representation of quantum states and operators.
@@
Although the literature on general-purpose TNs is extensive, there is considerable room for improvement regarding their application to quantum metrology. Among the existing approaches, Ref. [26] introduced a TN-based procedure to calculate quantum fidelity between connected subsystems of partially traced pure-state matrix product states. For mixed states, a more general method was proposed in Ref. [27], where both the quantum Fisher information (QFI) and the symmetric logarithmic derivative (SLD) are obtained through variational optimization. More recently, Ref. [28] presented an efficient framework for evaluating the quantum Cramér-Rao bound in continuously monitored systems subject to general Markovian or non-Markovian noise. In this paper, we introduce a novel approach suitable for evaluating the QFI and SLD from their integral forms. Although integral formulations are not new, our approach leverages tensor networks that enable scalable computations for large-scale systems. Importantly, the procedure can be efficiently implemented using imaginary-time evolution and basic tensor network algebra. This makes our approach directly applicable utilizing open source packages that include general MPS implementation, such as YASTN [29, 30], TeNPy [31] or ITensor [32]. In this paper, we analyze our method in the context of convergence time, the cost of numerical implementation, and numerical errors.
@@ I Preliminaries
The formal solution presented in Eq. (4) applies to the asymptotically stable Lyapunov equation and is equivalent to the pseudoinverse. The integral form remains valid when the density matrix is not full-rank, although additional regularization is necessary [34, 35, 40]. E.g. small admixture of the maximally mixed state $(1-\nu)\rho+\nu\frac{1}{\dim\rho}\mathbbm{1}$ where in the end we take $\nu\rightarrow 0$ [34]. An extended discussion of the SLD can be found in App. A. In this work, we restrict our attention to the cases where the SLD is well defined by the definition in Eq. (2) and the integral solution in Eq. (4) is valid.
@@ II Quantum metrology
Figure 1: Quantum metrology. The generally entangled input probe (white) is prepared in arbitrary quantum protocol, e.g., thermalization. Then, the probe is exposed to the encoding with the parameter ${\theta}$ by the local encoding unitary $\mathcal{U}_{\theta}$ (blue). Later, the probe is measured according to the optimal measurement in Eq. (2) (red). The measurement data are analyzed with statistical processing (yellow) to estimate the parameter value ${\theta}_{\rm est}$ and its variance $\delta^{2}{\theta}$.
Quantum metrology aims to leverage entanglement and superposition in order to improve the precision of the estimation of physical parameters. A typical estimation protocol follows a preparation-encoding-measurement sequence as visualized in Fig. 1: the probe is first prepared, then exposed to the signal, and later measured to analyze the data and estimate the signal amplitude. In the case of unitary encoding, the probe undergoes the following transformation
${{\rho}_{{\theta}}}=e^{-{\iota}{\theta}A}{\rho}_{0}\,e^{{\iota}{\theta}A},$ (6)
where ${\rho}_{0}$ is the state of the prepared probe, ${\theta}$ is the encoded parameter (often the parameter will be the accumulated phase, i.e. ${\theta}=\Theta t$ where $\Theta$ is the encoded parameter and $t$ is an interrogation time), and $A$ is the encoded operator. After encoding, the probe is measured and the parameter ${\theta}$ is estimated from the measurement data. The ultimate estimation precision is formalized by the celebrated Cramér-Rao bound [6, 2, 8, 11, 3]
@@
where $\delta^{2}{\theta}$ is a variance of the estimated parameter ${\theta}$ and ${\mathcal{F}_{\theta}}$ is the quantum Fisher information (QFI). The bound is asymptotically saturated in the limit of a large number of measurements through an optimal quantum measurement strategy which is encoded by the SLD in Eq. (2). Precisely, the optimal projective measurement is given by the eigenvectors of the SLD operator. For a specific measurement scheme, the bound in Eq. (7) takes the same form, but the QFI is replaced by the (classical) Fisher information calculated for the probability density function of measurement outcomes [11].
@@ IV Convergence
A general bound on the convergence error is difficult to calculate. Here, we present bounds on the error derived for unitary quantum metrology in Eq. (6) where the state derivative reads
@@ V.2 Tensor network implementation
Tensor networks provide a powerful framework for encoding quantum states and operators. For one-dimensional tensors and open boundary conditions of the model in Eq. (21) we choose an ansatz of matrix product operator (MPO) [45, 21]. In this representation, operators are decomposed into a series of rank-$4$ tensors, where two indices correspond to local bra and ket dimensions, while the others are so-called bond dimensions that encode correlations between local variables. The tensors form a one-dimensional network with a trivial dimension for the terminal bond dimension. The compression of the MPO relies on the Schmidt decomposition in subsequent bipartitions. The decomposition splits the operator into the sum $K=\sum_{j}\kappa_{j}L_{j}\otimes R_{j}$, where $\kappa_{j}$ are the Schmidt values and $L_{j}$ and $R_{j}$ are associated Schmidt left and right operators, respectively, such that ${\rm tr}[L_{i}L^{\dagger}_{j}]={\rm tr}[R_{i}R^{\dagger}_{j}]=\delta_{ij}$. To reduce the bond dimension, we truncate the Schmidt decomposition at the Schmidt value cutoff ${\rm SVtol.}$ such that $\kappa_{j}\geq{\rm SVtol.}$. Alternatively, we keep the maximal maximal bond dimension $D_{K}$ which fixes a number of significant contributions that we keep.
@@ V.5 Numerical errors
One of the sources of error is the accuracy with which the integrand is evaluated. In this paper, the integrand is propagated using the $2$nd-order TDVP algorithm for which the error scales as $\mathcal{O}({\Delta{x}}^{2})$, where ${\Delta{x}}$ is a discrete step for propagation of the integrand. To improve the accuracy of this step, we can apply higher-order TDVP procedures, however, this comes at increased computational cost. The second type of error is an accuracy of the numerical integration which depends on the integral discretization. For this step, we can employ standard numerical integration schemes, such as a quadrature rule, trapezoid rule, or higher-order Runge-Kutta (RK) methods [41]. The error of the RK-$k$ approach scales as $\mathcal{O}({\Delta{x}}^{k})$ where ${\Delta{x}}$ is a discrete integration step. For QFI-focused calculations, the integral can be evaluated in post-processing by taking the integrand as the input. This allows for flexible application of integration methods using fitting or interpolation procedures. In contrast to QFI, the SLD operator cannot be fitted externally and the integration has to be applied using the sum rule. The integration by a sum can be optimized by adopting an adaptive approach, see App. H, where the integration step is dynamically adjusted based on variations of the QFI integrand. The approach allows to control the integration error under the fixed tolerance at the cost of monitoring the QFI integrand. The adaptive procedure naturally allows step sizes to grow with increasing ${x}$ as the integrand becomes flatter and flatter. However, directly coupling the step size for integration to the step size for integrand propagation imposes a practical limitation, since the latter must be sufficiently small for TDVP. This restriction can be easily lifted when the step sizes are decoupled, but it may not be a significant gain since the TDVP cost remains dominant.
@@ Appendix C Low temperature approximation
that are the sums of elements of the encoding operator. We upper bound the expression by noticing that $\frac{(x-y)^{2}}{x+y}e^{-(x+y){X}}\leq|x-y|e^{-|x-y|{X}}$ where we have $|x-y|>0$ and obtain the expression
@@ Appendix E Unitary encoding
Figure 6: Errors of integration schemes. Comparison of the two implementations for the unitary encoding. The panels show (a) the QFI integrand and (b) the smallest Schmidt value for the integrand. Data are complementary to Fig. 4 and are obtained for the QFI integrand ${\bar{\mathcal{F}_{\theta}}}({x})=2{\rm tr}[\dot{{\rho}_{{\theta}}}{\mathcal{B}_{\theta}}({x})]$ where the ${\mathcal{B}_{\theta}}({x})$ as in Sec. III (solid) or Sec. V.3 (dashed). Despite the error for ${\mathcal{A}_{\theta}}$-based integration, its error on the QFI integral is $0.48\%$ of the ${\mathcal{F}_{\theta}}(50)$ value.
@@ Appendix F Cost of tensor network implementation
Figure 7: Dynamical compression. Comparison of the computational time for two scenarios: a fixed truncation threshold of ${\rm SVtol.}=10^{-9}$ (solid) and a dynamic compression approach where ${\rm SVtol.}({x})=10^{-9}\frac{{\bar{\mathcal{F}_{\theta}}}(0)}{{\bar{\mathcal{F}_{\theta}}}({x})}$ (dashed). The dynamic compression method employs more aggressive truncation when the integrand amplitude is low, thereby reducing computational costs. Data obtained for $N=16$, ${g}=2J$, $\beta=4J^{-1}$, bond dimensions $D_{{\rho}_{{\theta}}}\leq 32$, $D_{\mathcal{B}_{\theta}}\leq 64$ and integration step ${\Delta{x}}=0.1$. Computer time obtained for $4$ CPU cores.
@@ Appendix H Adaptive integration step
Figure 8: Adaptive integration step. (a) A comparison between the integration with a fixed step ${\Delta{x}}=0.1$ (blue) and the adaptive scheme with ${\rm tol\,}=0.1$ (red) and ${\rm tol\,}=0.01$ (green). The bound on the integration error for the fixed integration step in Eq. (60) is primarily accumulated in the beginning of simulation where the integrand decays quickly (dashed blue line and shaded area). (b) Numerical cost of the simulation measured by the computational time. For the adaptive scheme, necessity to make smaller steps when the integrand changes quickly may increase the computational cost. Data obtained for $N=16$, ${g}=2J$, $\beta=4J^{-1}$, bond dimensions $D_{{\rho}_{{\theta}}}\leq 32$ and $D_{\mathcal{B}_{\theta}}\leq 64$. Computer time obtained for $4$ CPU cores.

### v2

@@ Abstract
Although the role of QFI and SLD is crucial, their evaluation is notoriously difficult because of their nonlinear mathematical form. Among various formulas, a simplified version of the QFI for an adiabatic perturbation in a quantum many-body system can be derived [13]. In general, calculating the QFI requires matrix inversions, which become significant numerical bottlenecks. As a result, considerable effort has been devoted to developing methods for evaluating or bounding the QFI, paralleling advances in quantum metrology. The conventional approach, which relies heavily on eigendecomposition, faces a critical limitation: the Hilbert space dimension grows exponentially with the number of subsystems, making computations intractable for large, composite systems. To address these challenges, alternative formulations that avoid explicit diagonalization have been proposed [14, 15], although they still require certain matrix inversions. Other practical approaches provide QFI bounds using auxiliary measures such as the truncated QFI [16, 17], purity loss [18, 19], and the Krylov shadow method [20]. Recent experimental advances have used noisy intermediate-scale quantum (NISQ) devices to estimate these bounds [16, 17, 21], enabling scalable implementation on physical quantum platforms. In recent decades, tensor network (TN) [22, 23] methods have been shown to provide efficient approximations for relevant classes of states [24, 25, 23], e.g., ground states of gapped Hamiltonians and thermal states of a local Hamiltonian [26, 27]. Among them, the one-dimensional ansatz of the matrix product state (MPS) (also known as the matrix product operator (MPO) for quantum operators) provides a compact representation of quantum states and operators.
@@
Although the literature on general-purpose TNs is extensive, there is considerable room for improvement regarding their application to quantum metrology. Among the existing approaches, Ref. [28] introduced a TN-based procedure to calculate quantum fidelity between connected subsystems of partially traced pure-state matrix product states. For mixed states, a more general method was proposed in Ref. [29], where both the quantum Fisher information (QFI) and the symmetric logarithmic derivative (SLD) are obtained through variational optimization. The original implementation of the approach was released as the open-source package TNQMetro [30] (followed by QMetro++ [31]). More recently, Ref. [32] presented an efficient framework for evaluating the quantum Cramér-Rao bound in continuously monitored systems subject to general Markovian or non-Markovian noise. In this paper, we introduce a novel approach suitable for evaluating the QFI and SLD from their integral forms. Although integral formulations are not new, our approach leverages tensor networks that enable scalable computations for large-scale systems. Importantly, the procedure can be efficiently implemented using imaginary-time evolution and basic tensor network algebra. This makes our approach directly applicable utilizing open source packages that include general MPS implementation, such as YASTN [33, 34], TeNPy [35] or ITensor [36]. In this paper, we analyze our method in the context of convergence time, the cost of numerical implementation, and numerical errors.
@@ I Preliminaries
The formal solution presented in Eq. (4) applies to the asymptotically stable Lyapunov equation and is equivalent to the pseudoinverse. The integral form remains valid when the density matrix is not full-rank, although additional regularization is necessary [38, 39, 44]. In particular, it can be shown that the QFI matrix of any state can be calculated as a limit of the QFI matrix of a full-rank state [38], e.g. ${\mathcal{F}_{\theta}}=\lim_{\nu\rightarrow 0}\,(1-\nu)\rho_{\theta}+\nu\mathbbm{1}$ and $\nu\in(0,1)$. An extended discussion of the SLD can be found in App. A. In this work, we restrict our attention to the cases where the SLD is well defined by the definition in Eq. (2) and the integral solution in Eq. (4) is valid.
@@ II Quantum metrology
Figure 1: Typical procedure for quantum metrology. The generally entangled input probe (white) is prepared in arbitrary quantum protocol. The probe is exposed to the signal generated by the unitary $U_{\theta}$ that encodes the parameter ${\theta}$ (blue). After the encoding the probe is measured with the positive operator-valued measure $\{\Pi_{j}\}$ (red). The measurement data are analyzed with statistical processing (yellow) to estimate the parameter value ${\theta}_{\rm est}$ and its variance $\delta^{2}{\theta}$.
Quantum metrology aims to leverage entanglement and superposition in order to improve the precision of the estimation of physical parameters. A typical estimation protocol follows a preparation-encoding-measurement sequence as visualized in Fig. 1: a generally entangled quantum probe is first prepared; then the probe is exposed to the signal generated by the unitary $U_{\theta}$ that encodes the parameter ${\theta}$; later the probe is measured with the positive operator-valued measure defined by the set $\{\Pi_{j}\}$; the measurement data are analyzed to calculate the estimated value of the parameter ${\theta}_{\rm est}$ and the error of the value given by the variance $\delta^{2}{\theta}$.
The bound on the estimation precision, given large measurement statistics, is fixed by the celebrated (quantum) Cramér-Rao bound [6, 2, 8, 11, 3]
@@
where $\delta^{2}{\theta}$ is the variance of the estimated parameter ${\theta}$ and ${\mathcal{F}_{\theta}}$ is the Fisher information (FI) calculated for the probability density function of measurement outcomes [11]. The quantum Fisher information (QFI) is defined as the maximum value of the Fisher information achievable across all possible measurement strategies, and it is also closely related to the Bures metric, as discussed in Sec. I. The highest estimation precision is achieved through the optimal measurement, which is determined by the SLD operator defined in Eq. (2[6, 2, 8, 3]. Specifically, optimal projective measurement corresponds to the eigenvectors of the SLD operator.
@@ IV Convergence
A general bound on the convergence error is difficult to calculate. In this section, we focus on the scenario for typical for quantum metrology as explained in Sec. II. While the protocol for quantum metrology can be generalized, we will focus on the unitary encoding where the probe undergoes the following transformation
${{\rho}_{{\theta}}}=e^{-{\iota}{\theta}A}{\rho}_{0}\,e^{{\iota}{\theta}A},$ (13)
where ${\rho}_{0}$ is the state of the prepared probe, ${\theta}$ is the encoded parameter (often the parameter will be the accumulated phase, i.e. ${\theta}=\Theta t$ where $\Theta$ is the encoded parameter and $t$ is an interrogation time), and $A$ is the encoded operator generating the encoding unitary $U_{\theta}=e^{-{\iota}{\theta}A}$. There, the state derivative over the parameter ${\theta}$ reads
@@ V.2 Tensor network implementation
Tensor networks provide a powerful framework for encoding quantum states and operators. For one-dimensional tensors and open boundary conditions of the model in Eq. (21) we choose an ansatz of matrix product operator (MPO) [48, 23]. In this representation, operators are decomposed into a series of rank-444 tensors, where two indices correspond to local bra and ket dimensions, while the others are so-called bond dimensions that encode correlations between local variables. The tensors form a one-dimensional network with a trivial dimension for the terminal bond dimension. The compression of the MPO relies on the Schmidt decomposition in subsequent bipartitions. The decomposition splits the operator into the sum $K=\sum_{j}\kappa_{j}L_{j}\otimes R_{j}$, where $\kappa_{j}$ are the Schmidt values and $L_{j}$ and $R_{j}$ are associated Schmidt left and right operators, respectively, such that ${\rm tr}[L_{i}L^{\dagger}_{j}]={\rm tr}[R_{i}R^{\dagger}_{j}]=\delta_{ij}$. To reduce the bond dimension, we truncate the Schmidt decomposition at the Schmidt value cutoff ${\rm SVtol.}$ such that $\kappa_{j}\geq{\rm SVtol.}$. Alternatively, we keep the maximal maximal bond dimension $D_{K}$ which fixes a number of significant contributions that we keep.
@@ V.5 Numerical errors
[…]
@@ Appendix C Low temperature approximation
[…]
@@ Appendix E Unitary encoding
[…]
@@ Appendix F Cost of tensor network implementation
Figure 7: Dynamical compression. Comparison of the computational time for two scenarios: a fixed truncation threshold of ${\rm SVtol.}=10^{-9}$ (solid) and a dynamic compression approach where ${\rm SVtol.}({x})=10^{-9}\frac{{\bar{\mathcal{F}_{\theta}}}(0)}{{\bar{\mathcal{F}_{\theta}}}({x})}$ (dashed). The dynamic compression method employs more aggressive truncation when the integrand amplitude is low, thereby reducing computational costs. Data obtained for $N=16$, ${g}=2J$, $\beta=4J^{-1}$, bond dimensions $D_{{\rho}_{{\theta}}}\leq 32$, $D_{\mathcal{B}_{\theta}}\leq 64$ and integration step ${\Delta{x}}=0.1$. Computer time obtained for 444 CPU cores.
@@ Appendix H Adaptive integration step
Figure 8: Adaptive integration step. (a) A comparison between the integration with a fixed step ${\Delta{x}}=0.1$ (blue) and the adaptive scheme with ${\rm tol\,}=0.1$ (red) and ${\rm tol\,}=0.01$ (green). The bound on the integration error for the fixed integration step in Eq. (60) is primarily accumulated in the beginning of simulation where the integrand decays quickly (dashed blue line and shaded area). (b) Numerical cost of the simulation measured by the computational time. For the adaptive scheme, necessity to make smaller steps when the integrand changes quickly may increase the computational cost. Data obtained for $N=16$, ${g}=2J$, $\beta=4J^{-1}$, bond dimensions $D_{{\rho}_{{\theta}}}\leq 32$ and $D_{\mathcal{B}_{\theta}}\leq 64$. Computer time obtained for 444 CPU cores.